Antuan Company set the following standard costs for one unit of its product.

Direct materials (4.0 Ibs. @ $4.00 per Ib.) $ 16.00
Direct labor (1.7 hrs. @ $10.00 per hr.) 17.00
Overhead (1.7 hrs. @ $18.50 per hr.) 31.45
Total standard cost $ 64.45


The predetermined overhead rate ($18.50 per direct labor hour) is based on an expected volume of 75% of the factory’s capacity of 20,000 units per month. Following are the company’s budgeted overhead costs per month at the 75% capacity level.

Overhead Budget (75% Capacity)
Variable overhead costs
Indirect materials $ 15,000
Indirect labor 75,000
Power

15,000

Repairs and maintenance 30,000
Total variable overhead costs $ 135,000
Fixed overhead costs
Depreciation—Building 24,000
Depreciation—Machinery 70,000
Taxes and insurance 18,000
Supervision 224,750
Total fixed overhead costs 336,750
Total overhead costs $ 471,750


The company incurred the following actual costs when it operated at 75% of capacity in October.

Direct materials (61,000 Ibs. @ $4.10 per lb.) $ 250,100
Direct labor (22,000 hrs. @ $10.40 per hr.) 228,800
Overhead costs
Indirect materials $ 41,100
Indirect labor 176,300
Power 17,250
Repairs and maintenance 34,500
Depreciation—Building 24,000
Depreciation—Machinery 94,500
Taxes and insurance 16,200
Supervision 224,750 628,600
Total costs $ 1,107,500

Prepare flexible overhead budgets for October showing the amounts of each variable and fixed cost at the 65%, 75%, and 85% capacity levels and classify all items listed in the fixed budget as variable or fixed.

ANTUAN COMPANY
Flexible Overhead Budgets
For Month Ended October 31
Flexible Budget Flexible Budget for
Variable Amount per Unit Total Fixed Cost 65% of capacity 75% of capacity 85% of capacity
Sales (in units)
Variable overhead costs
$0.00 0 0 0
Fixed overhead costs
0 0 0 0
Total overhead costs
Answer #1
ANTUAN COMPANY
Flexible Overhead Budgets
For Month Ended October 31
Flexible Budget Flexible Budget for
Variable Amount per Unit Total Fixed Cost 65% of capacity 75% of capacity 85% of capacity
Sales (in units) (20000*65%)= 13000 (20000*75%)= 15000 (20000*85%)= 17000
Variable overhead costs
Indirect materials (15000/15000)= 1.00 (13000*$1.00)= 13000 (15000*$1.00)= 15000 (17000*$1.00)= 17000
Indirect labor (75000/15000)= 5.00 (13000*$5.00)= 65000 (15000*$5.00)= 75000 (17000*$5.00)= 85000
Power (15000/15000)= 1.00 (13000*$1.00)= 13000 (15000*$1.00)= 15000 (17000*$1.00)= 17000
Repairs and maintenance (30000/15000)= 2.00 (13000*$2.00)= 26000 (15000*$2.00)= 30000 (17000*$2.00)= 34000
Total variable overhead costs $9.00 $117000 $135000 $153000
Fixed overhead costs
Depreciation—Building 24000 24000 24000 24000
Depreciation—Machinery 70000 70000 70000 70000
Taxes and insurance 18000 18000 18000 18000
Supervision 224750 224750 224750 224750
Total fixed overhead costs $336750 $336750 $336750 $336750
Total overhead costs $453750 $471750 $489750

NOTE:- For any problem regarding the answer please ask in the comment section.